OpenStudy (anonymous):

Graph (x + 12)2 + (y + 3)2 = 100 by hand.

OpenStudy (anonymous):

www.analyzemath.com/high_school_math/grade_12/algebra.html gr. 12 right!!

OpenStudy (anonymous):

this is a linear function. so the graph is just a line. First let's rearrange it to point slope form given by, y=mx+b where the coefficient of x, m, is the slope and b is the y coordinate of the y-intercept... (x + 12)2 + (y + 3)2 = 100 ((x + 12)2 + (y + 3)2)/2 = 100/2 x+12+y+3=50 y=-x+50-12-3 y=-x+35 slope is -1 and the y intercept is (0,35) and setting y=0, will come up with the x-intercept also which is (35,0) also.. so the graph will be